OpenStudy (anonymous):

The first term of a geometric progression is 16 and the last term is 11664. If the sum of all the terms is 17488, determine the number of geometric means between 16 and 11664.

OpenStudy (anonymous):

very easy use the formula for g.p. and sum of g.p. solve for r and then a

OpenStudy (anonymous):

How do you solve for r?

OpenStudy (anonymous):

Let a be the first number, so a=16. The last number is \[ 16 r^k =11664 \\ r^k= 729= 3^6\\ \] r=3 and k=6 So there are 7 terms You really do not need to have the sum as given.

OpenStudy (anonymous):

\[ Sum= \frac{a \left(r^{k+1}-1\right)}{r-1}=\frac{16 \left(3^7-1\right)}{3-1}=17488 \] So you do not need that sum to be given to solve the problem. In fact, you can find it.

OpenStudy (anonymous):

Here are the seven terms \[ {16, 48, 144, 432, 1296, 3888, 11664} \]